OpenStudy (anonymous):

help with trig equation. Sin(2x)= a. 2sin(x) b. 2sin(x)cos(x) c. cos (x)-sin (x) 2 2 d. sin(2) + sin(x) e. none of these

OpenStudy (anonymous):

I don't want the answer at all I want the process

OpenStudy (unklerhaukus):

Using the Sum & Difference Formula (for sine) \[\sin(u ± v) = \sin(u)\cos(v)± \cos(u)\sin(v)\] for your expression \[u=v=x\]

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

sin(2x)=sin(x+x)=

OpenStudy (anonymous):

you get\[2\sin x \cos x\]

OpenStudy (unklerhaukus):

yes that is correct

OpenStudy (anonymous):

ok, and I'm guessing it's not the same for cos, tan, sec, csc, or the inverses

OpenStudy (unklerhaukus):

The Sum & Difference Formulas for Cosine \[\cos(u ± v) = \cos(u)\cos(v)∓ \sin(u)\sin(v)\] and Tangent \[\tan(u±v)=\frac{\tan(u)±\tan(v) }{1∓\tan(u)\tan(v)}\]

OpenStudy (anonymous):

b. 2sin(x)cos(x)

OpenStudy (anonymous):

Ok, thank you very much, what level trig is this from? I'm trying to refresh myself on everything I took in high school as I am attending university next september but I've been out of school for 5 years so some of it has been completely lost, mostly the trig though, a lot of the calc came back prettyquick and algebra was never really lost

OpenStudy (anonymous):

or rather the basic algebra was never lost, some of the more advanced algebra is taking a little bit of time to remember

OpenStudy (unklerhaukus):

I don't attempt to remember these formulas I look them up when i need them

OpenStudy (anonymous):

thank you very much :) btw does \[(\sqrt{x}+\sqrt{y})^{2}=x+y\] or \[(\sqrt{x}+\sqrt{y})^{2}=x+2\sqrt{x}\sqrt{y}+y\]

OpenStudy (ranga):

The second one because\[\Large (a + b)^2 = a^2 + 2ab + b^2\]

OpenStudy (anonymous):

ok thought so just wanted to make sure I wasn't forgetting yet another rule. So many I've forgotten and need to remember :/

OpenStudy (ranga):

It will all come back pretty fast though.

OpenStudy (unklerhaukus):